OpenStudy (anonymous):

solve algebraically 8cos(2x-5)=3 general solution in radians 5.2sin (45degrees(x+8degrees))-1=-3 general solution in degrees

OpenStudy (campbell_st):

for the 1st equation divide booth sides by 8 \[\cos(2x - 5) = \frac{3}{8}\] so then \[2x - 5 = \cos^{-1}(\frac{3}{8})\] find the angle in radians, then add 5 and divide by 2 to find x